OpenStudy (anonymous):

one molecule of a compound contains six carbon atoms, 2*10^-23g of hydrogen and 16* 10^-23 of other atoms the gram molecular mass of compound is ??

OpenStudy (jfraser):

cant be done with the information you've supplied

OpenStudy (anonymous):

it is the question in my book

OpenStudy (anonymous):

options are 180g 174g 138g 114g

OpenStudy (anonymous):

the answer is 180g but i dont know to solve it

OpenStudy (anonymous):

which board are you in?

OpenStudy (anonymous):

cbse

OpenStudy (anonymous):

oh.I don't remember solving such a question

OpenStudy (anonymous):

it is an extra question not from the text book

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

okay its like this...i ve worked it out and i came up with this solution... Now they have given the mass of Hydrogen and *other atoms*= when u add both and divide by the Avagrado constant, You get the mass (my bad i forgot it :P) \[[2\times 10^{-23} + 16 \times 10^{-23} = 1.8 \times 10^{-22} \times 6.022 \times 10^{23}\] = Mass of the (Hydrogen + other substance) =108 (approximately) ADD it with 6 carbon atoms ------> 180 :)

OpenStudy (anonymous):

not divide it should be multiply

OpenStudy (anonymous):

i did nt understand can u make it clear

OpenStudy (anonymous):

sorry now i understood.............